Find two consecutive even integers whose product is 24 more than 8 times the larger integer
Answer by stanbon(75887)

Find two consecutive even integers whose product is 24 more than 8 times the larger integer
---
1st: 2x
2nd: 2x+2
---------
Equation:
2x(2x+2) = 8(2x+2) + 24
----
4x^2+4x = 16x+16 + 24
----
4x^2 - 12x - 40 = 0
---
x^2 -3x - 10 = 0
(x-5)(x+2) = 0
-------------------
If x = 5
2x = 10
and 2x+2 = 12
---
If x = -2
2x = -4
2x+2 = -2
----
=================
Cheers,
Stan H.
===============